Find the general solution of the differential equation .
y²/2 + 2yexp(t) + (y+exp(t))dy/dt = 0
1
Expert's answer
2019-08-22T10:28:52-0400

"y^2\/2+2ye^t+(y+e^t)y'=0"

Multiply by 2

"y^2+4ye^t+2(y+e^t)y'=0"

"y^2+2ye^t+e^{2t}+2(y+e^t)y'+2ye^t+2e^{2t}-3e^{2t}=0"

"(y+e^t)^2+2(y+e^t)y'+2(y+e^t)e^t-3e^{2t}=0"

"(y+e^t)^2+2(y+e^t)(y'+e^t)=3e^{2t}"

Notice that

"((y+e^t)^2)'=2(y+e^t)(y'+e^t)"

Let "z=(y+e^t)^2"

then

"z+z'=3e^{2t}"

Using an integrating factor

"u(t)=e^{\int1dt}=e^t"

General solution is

"z=\frac{\intop e^t3e^{2t}dt+C}{e^t}=\frac{e^{3t}+C}{e^t}"

"(y+e^t)^2=\frac{e^{3t}+C}{e^t}"

"y+e^t=\pm \sqrt{\frac{e^{3t}+C}{e^t}}"

"y=\pm \sqrt{\frac{e^{3t}+C}{e^t}}-e^t"

Answer: "y=\pm \sqrt{\frac{e^{3t}+C}{e^t}}-e^t"